### v1

# Path-dependent Hamilton–Jacobi equations with $u$-dependence and time-measurable Hamiltonians

Abstract.
We establish existence and uniqueness of minimax solutions for a fairly general class of path-dependent Hamilton–Jacobi equations. In particular, the relevant Hamiltonians can contain the solution and they only need to be measurable with respect to time. We apply our results to optimal control problems of (delay) functional differential equations with cost functionals that have discount factors and with time-measurable data. Our main results are also crucial for our companion paper Bandini and Keller (2024), where non-local path-dependent Hamilton–Jacobi–Bellman equations associated to the stochastic optimal control of non-Markovian piecewise deterministic processes are studied.
a University of Bologna, Department of Mathematics, Piazza di Porta San Donato 5, 40126 Bologna (Italy).
b University of Central Florida, Department of Mathematics, 4393 Andromeda Loop N Orlando, FL 32816 (USA).
1 E-mail: elena.bandini7@unibo.it.
2 E-mail: christian.keller@ucf.edu.
The research of the first named author was partially supported by the 2018 GNAMPA-INdAM project Controllo ottimo stocastico con osservazione parziale: metodo di randomizzazione ed equazioni di Hamilton-Jacobi-Bellman sullo spazio di Wasserstein and the 2024 GNAMPA-INdAM Problemi di controllo ottimo in dimensione infinita. The research of the second named author was supported in part by NSF-grant DMS-2106077.
Keywords: Path-dependent Hamilton–Jacobi equations; time-measurable Hamiltonians; minimax solutions; comparison principle; optimal control.
AMS 2020: 35F21, 49L25, 34K35
1. Introduction
We study path-dependent Hamilton–Jacobi equations of the form
(1.1) $\partial_{t}u(t,x)+H(t,x,u(t,x),\partial_{x}u(t,x))=0,\quad(t,x)\in[0,T)\times D([0,T],{\mathbb{R}}^{d}),$
with time-measurable Hamiltonian, i.e., the mappings $t\mapsto H(t,x,y,z)$ only need to be Borel-measurable. The path space $D([0,T],{\mathbb{R}}^{d})$ in (1.1) is the set of all right-continuous functions from $[0,T]$ to ${\mathbb{R}}^{d}$ that have left limits. Important special cases of (1.1) are Isaacs equations associated to differential games with history-dependent data and Hamilton–Jacobi–Bellman (HJB) equations associated to optimal controls problems involving (delay) functional differential equations. There exists already a large body of literature for non-smooth solutions of those equations with continuous Hamiltonians (see [GLP21AMO] and the references therein). Typically, $C([0,T],{\mathbb{R}}^{d})$ is used as path space. But this difference is not material.
The two most common notions of non-smooth solutions for path-dependent Hamilton–Jacobi equations are viscosity solutions and minimax solutions. In fact, for a large class of path-dependent equations but with continuous (!) Hamiltonians, these two notions turn out to be equivalent (see [GP23JFA]).111 Note that we do not pursue here a generalization of those equivalence results in our setting.
Here, we work with minimax solutions. This helps a lot in our setting, where $H$ in (1.1) is time-measurable, as no change of notion is needed compared to the case of $H$ being continuous. Furthermore, we also extend the theory of minimax solutions for path-dependent equations in [Lukoyanov03] to the case of $u$-dependent Hamiltonians and thereby provide a counterpart to the theory of minimax solutions for (non-path-dependent) Hamilton–Jacobi equations with $u$-dependent Hamiltonians in [Subbotin] (see also the references therein for earlier works but with more restrictive assumptions).
Viscosity solutions still play a useful role. They are utilized for the proof of our comparison principle.222 We use the same methodology for establishing existence and uniqueness for minimax solutions with the help of viscosity solution techniques as in [BK18JFA]: The “five-step-scheme” described on p. 2103 therein to be more precise.
Finally note that we choose $D([0,T],{\mathbb{R}}^{d})$ as path space and we work with relatively weak assumptions for $H$ (in particular, time-measurability) because doing so is crucial for establishing well-posedness of non-local path-dependent partial differential equations in our companion paper [BK2] (see section 1.2 therein for details).
1.1. Further related literature
Results for viscosity solutions for (non-path-dependent) equations with only time-measurable Hamiltonians were first published in [Ishii85] (this work had also some influence for the proof of our comparison principle in section 5). Further related early works are [LionsPerthame87] and [BarronJensen87].
Minimax solutions for (non-path-dependent) equations with time-measurable Hamiltonians were studied in [SonEtAl95, SonEtAl97, minimaxBookTimeMeas, VanEtAl95]. Note that these works unlike ours require positive homogeneity of the Hamiltonian with respect to the gradient.
In [Qiu22SPA_RHJB], a quite similar optimal control problem (compared to ours in section 7) is studied. Besides being time-measurable, the coefficients in [Qiu22SPA_RHJB] can even be random. Correspondingly, the controller minimizes an expected cost functional. However, in order to prove uniqueness for the associated HJB equation, time-continuity is required in contrast to our work. Moreover, we cover more general Hamilton–Jacobi equations besides HJB equations.
1.2. A difficulty and its resolution
To prove a comparison principle for viscosity solutions, it is typical to consider the doubled equation
(1.2) $\partial_{t}w+H(t,x,u,\partial_{x}w)-H(t,\tilde{x},v,-\partial_{\tilde{x}}w)=0$
[…]
1.3. Organization of the rest of the paper
Section 2 introduces the setting (path space with topology) and some notation. In section 3, we give meaning to the derivatives $\partial_{t}u$ and $\partial_{x}u$ in (1.1). Section 4 contains the definition of minimax solution to terminal-value problems involving (1.1) as well as standing assumptions for our data such as the Hamiltonian $H$. In section 5, a comparison principle for (1.1) is established. In section 6, we establish a general existence result via Perron’s method. Finally, in section 7, we consider an optimal control problem for (delay) functional differential equations with time-measurable data and we show that the value function is the unique minimax solution to the associated HJB equation.
2. Setting and notation
Let $\Omega:=D([0,T],{\mathbb{R}}^{d})$. We equip $\Omega$ with the supremum norm $\lVert{\cdot}\rVert_{\infty}$ and $[0,T]\times\Omega$ with a pseudo-metric $\mathbf{d}_{\infty}$ defined by
$\mathbf{d}_{\infty}((t,x),(s,\tilde{x})):=\left|{t-s}\right|+\sup_{0\leq r\leq T}\left|{x(r\wedge t)-\tilde{x}(r\wedge s)}\right|.$
Given a set $S\subset[0,T]$, we write ${\mathbf{1}}_{S}$ for the corresponding indicator function, i.e., ${\mathbf{1}}_{S}(t)=1$ if $t\in S$ and ${\mathbf{1}}_{S}(t)=0$ if $t\in[0,T]\setminus S$.
Given topological spaces $E$ and $F$, we denote by $C(E,F)$ the set of all continuous functions from $E$ to $F$. In case $F={\mathbb{R}}$, we just write $C(E)$. Similarly, we denote by $\mathrm{USC}(E)$ the set of all upper semicontinuous (u.s.c.) functions from $E$ to ${\mathbb{R}}$ and by $\mathrm{LSC}(E)$ the set of all lower semicontinuous (l.s.c.) functions from $E$ to ${\mathbb{R}}$.
We write $a\cdot b=(a,b)$ for the inner product of two vectors $a$ and $b$ in ${\mathbb{R}}^{d}$.
Given $L\geq 0$, $(s_{0},x_{0})\in[0,T)\times\Omega$, define
$\mathcal{X}^{L}(s_{0},x_{0})$ :=$:=\{x\in\Omega:\,x=x_{0}\text{ on $[0,s_{0}]$, $x|_{[s_{0},T]}$ is absolutely continuous with}$
$\qquad\qquad\left|{x^{\prime}(t)}\right|\leq L(1+\sup_{s\leq t}\left|{x(s)}\right|)\text{ a.e.~{}on $(t_{0},T)$}\}.$
Those sets of “Lipschitz-like” paths are very important. In particular, thanks to their useful properties listed in the following remarks, they are helpful insofar they circumvent difficulties coming from the lack of local compactness of $\Omega$.
Remark 2.1.
The sets $\mathcal{X}^{L}(s_{0},x_{0})$ are compact in $(\Omega,\lVert{\cdot}\rVert_{\infty})$ (see, e.g., Proposition 4.1 in [Lukoyanov03] or Proposition 2.10 in [BK18JFA]).
Remark 2.2.
Let $L\geq 0$ and $(t_{n},x_{n})\to(t_{0},x_{0})$ in $[0,T]\times\Omega$ as $n\to\infty$. Then every sequence $(\tilde{x}_{n})_{n}$ with $\tilde{x}_{n}\in\mathcal{X}^{L}(t_{n},x_{n})$, $n\in{\mathbb{N}}$, has a subsequence that converges to some $\tilde{x}_{0}\in\mathcal{X}^{L}(t_{0},x_{0})$ (cf. Proposition 4.2 in [Lukoyanov03] and Proposition 2.12 in [BK18JFA]). For a detailed proof, follow the approach of Lemma 1 on page 87 in [FilippovBook].
3. Path derivatives
Our path derivatives are due to A. V. Kim (see [KimBook] for a detailed exposition).
Definition 3.1.
We write $\varphi\in C^{1,1,1}([0,T]\times\Omega\times\Omega)$ if $\varphi\in C([0,T]\times\Omega\times\Omega)$ and there are functions $\partial_{t}\varphi\in C([0,T]\times\Omega\times\Omega)$ and $\partial_{x}\varphi$, $\partial_{\tilde{x}}\varphi\in C([0,T]\times\Omega\times\Omega,{\mathbb{R}}^{d})$, called path derivatives of $\varphi$, such that, for every $(t_{0},x_{0},\tilde{x}_{0})\in[0,T)\times\Omega\times\Omega$, every pair $(x,\tilde{x})\in C([0,T]\times\Omega)$ with $(x,\tilde{x})|_{[0,t_{0}]}=(x_{0},\tilde{x}_{0})|_{[0,t_{0}]}$ and $(x,\tilde{x})|_{[t_{0},T]}$ being Lipschitz continuous, and every $t\in(t_{0},T]$, we have
$\varphi(t,x,\tilde{x})-\varphi(t_{0},x_{0},\tilde{x}_{0})$
$\qquad=\int_{t_{0}}^{t}\left[\partial_{t}\varphi(s,x,\tilde{x})+x^{\prime}(s)\cdot\partial_{x}\varphi(s,x,\tilde{x})+\tilde{x}^{\prime}(s)\cdot\partial_{\tilde{x}}\varphi(s,x,\tilde{x})\right]\,{\mathrm{d}}s.$
Remark 3.2.
The path derivatives of any function in $C^{1,1,1}([0,T]\times\Omega\times\Omega)$ are uniquely determined (see, e.g., Remark 2.17 in [BK18JFA]).
4. Minimax solutions
Fix functions $H:[0,T]\times\Omega\times{\mathbb{R}}\times{\mathbb{R}}^{d}\to{\mathbb{R}}$ and $h:\Omega\to{\mathbb{R}}$. We consider the terminal-value problem
(4.1) $\begin{split}-\partial_{t}u-H(t,x,u,\partial_{x}u)&=0,\quad(t,x)\in[0,T)\times\Omega,\\ u(T,x)&=h(x),\quad x\in\Omega.\end{split}$
The following two assumptions are always in force.
Assumption 4.1.
The function $h$ is continuous.
Assumption 4.2.
Suppose that $H$ satisfies the following conditions.
(i) For a.e. $t\in(0,T)$, the function $(x,y,z)\mapsto H(t,x,y,z)$, $\Omega\times{\mathbb{R}}\times{\mathbb{R}}^{d}\to{\mathbb{R}}$, is continuous.
(ii) For every $(x,y,z)\in\Omega\times{\mathbb{R}}\times{\mathbb{R}}^{d}$, the function $t\mapsto H(t,x,y,z)$, $[0,T]\to{\mathbb{R}}$, is Borel measurable.
(iii) There is a constant $L_{H}\geq 0$ such that, for a.e. $t\in(0,T)$, every $x\in\Omega$, $y\in{\mathbb{R}}$, $z$, $\tilde{z}\in{\mathbb{R}}^{d}$,
$\left|{H(t,x,y,z)-H(t,x,y,\tilde{z})}\right|\leq L_{H}(1+\sup_{s\leq t}\left|{x(s)}\right|)\left|{z-\tilde{z}}\right|.$
(iv) For every $L\geq 0$, there exists a constant $M_{L}\geq 0$ such that, for every $(t_{0},x_{0})\in[0,T)\times\Omega$, $x$, $\tilde{x}\in\mathcal{X}^{L}(t_{0},x_{0})$, $y\in{\mathbb{R}}$, $z\in{\mathbb{R}}^{d}$, and a.e. $t\in(t_{0},T)$,
$\left|{H(t,x,y,z)-H(t,\tilde{x},y,z)}\right|$ $\leq M_{L}(1{\color[rgb]{0,0,0}+\left|{y}\right|}+\left|{z}\right|)\,\sup_{s\leq t}\left|{x(s)-\tilde{x}(s)}\right|.$
(v) For a.e. $t\in(0,T)$ and every $(x,z)\in\Omega\times{\mathbb{R}}^{d}$, the function $y\mapsto H(t,x,y,z)$, ${\mathbb{R}}\to{\mathbb{R}}$, is non-increasing.
(vi) There is a constant $C_{H}\geq 0$ such that, for a.e. $t\in(0,T)$ and all $(x,y)\in\Omega\times{\mathbb{R}}$,
$\left|{H(t,x,y,0)}\right|\leq C_{H}(1+\sup_{s\leq t}\left|{x(s)}\right|+\left|{y}\right|).$
Next, we introduce sets of paths needed in our definition of minimax solution.555 Note that without the $u$-dependence of $H$, the situation would be much easier. Only the sets $\mathcal{X}^{L}(s_{0},x_{0})$ (and not $\mathcal{Y}^{L}(s_{0},x_{0},y_{0},z)$) would then be needed (see, e.g., [BK18JFA, section 1.3]).
Given $L\geq 0$, $s_{0}\in[0,T)$, $x_{0}\in\Omega$, $y_{0}\in{\mathbb{R}}$, and $z\in{\mathbb{R}}^{d}$, define
$\mathcal{Y}^{L}(s_{0},x_{0},y_{0},z)$ :=$:=\{(x,y)\in\mathcal{X}^{L}(s_{0},x_{0})\times C([t_{0},T]):$
$\qquad\qquad y(t)=y_{0}+\int_{s_{0}}^{t}\left[(x^{\prime}(s),z)-H(s,x,y(s),z)\right]\,{\mathrm{d}}s\text{ on $[t_{0},T]$}\}.$
Remark 4.3.
Thanks to Assumption 4.2 (i) and (vi), the sets $\mathcal{Y}^{L}(s_{0},x_{0},y_{0},z)$ are non-empty (cf. Proposition 2.13 (i) in [BK18JFA]) and compact in $(\Omega\times C([t_{0},T]),\lVert{\cdot}\rVert_{\infty})$ (cf.  Remark 2.1). Moreover, using additionally Assumption 4.2 (iii), one can show that also the intersections $\mathcal{Y}^{L_{H}}(s_{0},x_{0},y_{0},z)\cap\mathcal{Y}^{L_{H}}(s_{0},x_{0},y_{0},\tilde{z})$, $z\neq\tilde{z}$, are non-empty (see, e.g., [Subbotin, pages 73-74] or [BK18JFA, pages 2124-2125]).
Remark 4.4.
Let $L\geq 0$, $z\in{\mathbb{R}}^{d}$, and $(t_{n},x_{n},y_{n})\to(t_{0},x_{0},y_{0})$ in $[0,T]\times\Omega\times{\mathbb{R}}$ as $n\to\infty$. Then every sequence $(\tilde{x}_{n},\tilde{y}_{n})_{n}$ in $\Omega\times D([0,T])$ with $(\tilde{x}_{n},\tilde{y}_{n}|_{[t_{n},T]})\in\mathcal{Y}^{L}(t_{n},x_{n},y_{n},z)$, $n\in{\mathbb{N}}$, has a subsequence that converges to some $(\tilde{x}_{0},\tilde{y}_{0})$ in $(\Omega\times D([0,T]),\lVert{\cdot}\rVert_{\infty})$ with $(\tilde{x}_{0},\tilde{y}_{0}|_{[t_{0},T]})\in\mathcal{Y}^{L}(t_{0},x_{0},y_{0},z)$. This follows from Remark 2.2 and an appropriate adaption of the proof of Lemma 5 on page 8 in [FilippovBook] to our setting.
Definition 4.5.
Let $L\geq 0$ and $u:[0,T]\times\Omega\to{\mathbb{R}}$.
(i) $u$ is a minimax $L$-supersolution of (4.1) if $u\in\mathrm{LSC}([0,T]\times\Omega)$, if $u(T,\cdot)\geq h$ on $\Omega$, and if, for every $(s_{0},x_{0},z)\in[0,T)\times\Omega\times{\mathbb{R}}^{d}$, and every $y_{0}\geq u(s_{0},x_{0})$, there exists an $(x,y)\in\mathcal{Y}^{L}(s_{0},x_{0},y_{0},z)$ such that $y(t)\geq u(t,x)$ for each $t\in[s_{0},T]$.
(ii) $u$ is a minimax $L$-subsolution of (4.1) if $u\in\mathrm{USC}([0,T]\times\Omega)$, if $u(T,\cdot)\leq h$ on $\Omega$, and if, for every $(s_{0},x_{0},z)\in[0,T)\times\Omega\times{\mathbb{R}}^{d}$, and every $y_{0}\leq u(s_{0},x_{0})$, there exists an $(x,y)\in\mathcal{Y}^{L}(s_{0},x_{0},y_{0},z)$ such that $y(t)\leq u(t,x)$ for each $t\in[s_{0},T]$.
(iii) $u$ is a minimax $L$-solution of (4.1) if it is both a minimax $L$-supersolution and a minimax $L$-subsolution of (4.1).
Remark 4.6.
Notice that classical solutions are minimax solutions. We sketch some of the arguments (adapted from [Subbotin, section 2.4 ]). Assume that $H$ is continuous and that $u$ is a classical solution of (4.1), i.e., $u\in C^{1,1}([0,T]\times\Omega)$ (this space is an obvious modification666 Just assume that $(t,x,\tilde{x})\mapsto u(t,x)\in C^{1,1,1}([0,T]\times\Omega)$. of Definition 3.1) and, for every $(t,x)\in[0,T)\times\Omega$,
$-\partial_{t}u(t,x)-H(t,x,u(t,x),\partial_{x}u(t,x))=0.$
We will show that $u$ is a minimax $L_{H}$-supersolution of (4.1) with the slight modification that instead of $y_{0}\geq u(s_{0},x_{0})$ in Definition 4.5 (i), we only require to consider the case $y_{0}=u(s_{0},x_{0})$. To this end, fix $(s_{0},x_{0},z)\in[0,T)\times\Omega\times{\mathbb{R}}^{d}$ and let $y_{0}=u(s_{0},x_{0})$. Let $x$ be a solution of
$x^{\prime}(t)=\begin{cases}\frac{H(t,x,u(t,x),\partial_{x}u(t,x))-H(t,x,u(t,x),z)}{\left|{\partial_{x}u(t,x)-z}\right|^{2}}\cdot(\partial_{x}u(t,x)-z)&\text{if $u(t,x)\neq z$,}\\ 0&\text{if $u(t,x)=z$}\end{cases}.$
a.e. on $(s_{0},T)$ with initial condition $x(t)=x_{0}(t)$ for every $t\in[0,s_{0}]$. Note that $x\in\mathcal{X}^{(L_{H})}(s_{0},x_{0})$ according to Assumption 4.2. Let $y$ be a solution of
$y^{\prime}(t)=(x^{\prime}(t),z)-H(t,x,u(t,x),z)\quad\text{a.e.~{}on $(s_{0},T)$}$
with initial condition $y(s_{0})=y_{0}$. Since
$\frac{d}{dt}u(t,x)$ $=\frac{d}{dt}\int_{s_{0}}^{t}\partial_{t}u(s,x)+(x^{\prime}(s),\partial_{x}u(s,x))\,ds$
$=-H(t,x,u(t,x),\partial_{x}u(t,x))+(x^{\prime}(t),\partial_{x}u(t,x)-z)+(x^{\prime}(t),z)$
$=-H(t,x,u(t,x),z)+(x^{\prime}(t),z)=y^{\prime}(t),$
we have $(x,y)\in\mathcal{Y}^{(L_{H})}(s_{0},x_{0},y_{0},z)$ and $y(t)=u(t,x)$ for all $t\in[s_{0},T]$. Hence, $u$ is a minimax $L_{H}$-supersolution of (4.1) (in the slightly modified sense specified at the beginning of this remark).
The next lemma provides an equivalent criterion for a function to be a minimax supersolution. A corresponding statement holds for minimax subsolutions.
Lemma 4.7.
A function $u:[0,T]\times\Omega\to{\mathbb{R}}$ is a minimax $L$-supersolution of (4.1) if and only if $u\in\mathrm{LSC}([0,T]\times\Omega)$, $u(T,\cdot)\geq h$, and, for each $(s_{0},x_{0},z)\in[0,T)\times\Omega\times{\mathbb{R}}^{d}$, $y_{0}\geq u(s_{0},x_{0})$, and $t\in(s_{0},T]$, there is an $(x,y)\in\mathcal{Y}^{L}(s_{0},x_{0},y_{0},z)$ such that $y(t)\geq u(t,x)$.
Proof.
We only prove the non-trivial direction. We proceed along the lines of the proof of Lemma 3.6 in [BK18JFA]. Fix $(s_{0},x_{0},z)\in[0,T)\times\Omega\times{\mathbb{R}}^{d}$ and $y_{0}\geq u(s_{0},x_{0})$. Given is the following:
(4.2) $\begin{split}&\forall(s_{1},x_{1})\in[s_{0},T)\times\Omega:\,\forall y_{1}\geq u(s_{1},x_{1}):\,\forall t\in(s_{0},T]:\\ &\qquad\exists(x,y)\in\mathcal{Y}^{L}(s_{1},x_{1},y_{1},z):\,y(t)\geq u(t,x).\end{split}$
We have to establish the existence of a pair $(x,y)\in\mathcal{Y}^{L}(s_{0},x_{0},y_{0},z)$ independent from $t$ such that $y\geq u(\cdot,x)$ on $(s_{0},T)$. To this end, consider a sequence $(\pi^{m})_{m}$ of dyadic partitions of $[s_{0},T]$ with $\pi^{m}:s_{0}=t_{0}^{m}<t_{1}^{m}<\cdots<t^{m}_{n(m)}=T$, $m\in{\mathbb{N}}$, and $\sup_{i}(t^{m}_{i+1}-t^{m}_{i})\to 0$ as $m\to\infty$. By (4.2), there exists, for each $m\in{\mathbb{N}}$, a finite sequence $(x^{m}_{i},y^{m}_{i})_{i=1}^{n(m)}$ such that we have $(x^{m}_{1},y^{m}_{1})\in\mathcal{Y}^{L}(s_{0},x_{0},y_{0},z)$ with $y^{m}_{1}(t^{m}_{1})\geq u(t^{m}_{1}),x^{m}_{1})$ and, for each $i\in\{1,\ldots,n(m)-1\}$, we have $(x^{m}_{i+1}),y^{m}_{i+1})\in\mathcal{Y}^{L}(t^{m}_{i},x^{m}_{i},y^{m}_{i}(t^{m}_{i}),z)$ with $y^{m}_{i+1}(t^{m}_{i+1})\geq u(t^{m}_{i+1},x^{m}_{i+1})$. Using those sequences, we define pairs $(x^{m},y^{m})\in\mathcal{Y}^{L}(s_{0},x_{0},y_{0},z)$, $m\in{\mathbb{N}}$, by
$x^{m}(t)=\begin{cases}x_{0}&\text{if $0\leq t\leq s_{0}$,}\\ x^{m}_{i}(t)&\text{if $t^{m}_{i-1}<t\leq t^{m}_{i}$},\end{cases}\quad\text{and}\quad y^{m}(t)=\begin{cases}y_{0}&\text{if $t=s_{0}$,}\\ y^{m}_{i}(t)&\text{if $t^{m}_{i-1}<t\leq t^{m}_{i}$}.\end{cases}$
By compactness777 Note that Assumption 4.2 (vi) prevents a possible blow up of $y^{m}$ as $m\to\infty$. of $\mathcal{Y}^{L}(s_{0},x_{0},y_{0},z)$ (Remark 4.3), we can, without loss of generality, assume that $(x^{m},y^{m})_{m}$ converges uniformly to a pair $(x^{0},y^{0})\in\mathcal{Y}^{L}(s_{0},x_{0},y_{0},z)$. Let $t\in(s_{0},T]$. Then there is a sequence $(s^{m})_{m}$ in $(s_{0},T]$ with $s^{m}\in\pi^{m}$, $m\in{\mathbb{N}}$, that converges to $t$. Thus $y^{0}(t)=\lim_{m}y^{m}(s^{m})\geq\liminf_{m}u(s^{m},x^{m})\geq u(t,x^{0})$ thanks to the lower semi-continuity of $u$. This concludes the proof. ∎
5. Comparison principle
First, we introduce viscosity solutions for a suitable doubled equation (equation (5.2) below instead of the “naive” doubled equation (1.2)). Next, we establish connections between minimax solutions of (4.1) and those viscosity solutions. Finally, we prove a comparison principle for our doubled equation, which immediately leads to a comparison principle between minimax sub- and supersolutions of (4.1).
We start by defining spaces of test functions, which are needed for our notion of viscosity solutions: Given $L\geq 0$, $(s_{0},x_{0},\tilde{x}_{0})\in[0,T)\times\Omega\times\Omega$, and a function $w:[0,T]\times\Omega\times\Omega\to{\mathbb{R}}$, let
(5.1) $\begin{split}&\underline{\mathcal{A}}^{L}w(s_{0},x_{0},\tilde{x}_{0}):=\{\varphi\in C^{1,1,1}([s_{0},T]\times\Omega\times\Omega):\,\exists T_{0}\in(t_{0},T]:\\ &\qquad 0=(\varphi-w)(s_{0},x_{0},\tilde{x}_{0})=\inf\{(\varphi-w)(t,x,\tilde{x}):\\ &\qquad\qquad\qquad(t,x,\tilde{x})\in[s_{0},T_{0}]\in\mathcal{X}^{L}(s_{0},x_{0})\times\mathcal{X}^{L}(s_{0},\tilde{x}_{0})\}\}.\end{split}$
The next definition is vaguely inspired by [Ishii85].
Definition 5.1.
Fix $L\geq 0$ and a function $\Upsilon:[0,T]\times\Omega\to{\mathbb{R}}$. Let $M_{L}$ be the constant from Assumption 4.2 (iv). A function $w:[0,T]\times\Omega\times\Omega\to{\mathbb{R}}$ is an $L$-viscosity subsolution of
(5.2) $\begin{split}&\max\{w,0\}\cdot\Bigl{[}\partial_{t}w+M_{L}\left(1{\color[rgb]{0,0,0}+\left|{\Upsilon}\right|}+\left|{\partial_{x}w}\right|\right)\,\sup\nolimits_{t\leq s}\left|{x(t)-\tilde{x}(t)}\right|\\ &\qquad\qquad\qquad\qquad{\color[rgb]{0,0,0}+L_{H}\left(1+\sup\nolimits_{t\leq s}\left|{\tilde{x}(t)}\right|\right)\,\left|{\partial_{x}w+\partial_{\tilde{x}}w}\right|}\Bigr{]}=0\end{split}$
on $[0,T)\times\Omega\times\Omega$ with parameter $\Upsilon$ if $w$ is u.s.c. and, for every $(s,x,\tilde{x})\in[0,T)\times\Omega\times\Omega$ and every test function $\varphi\in\underline{\mathcal{A}}^{L}w(s,x,\tilde{x})$, we have
(5.3) $\begin{split}&\partial_{t}\varphi(s,x,\tilde{x})+M_{L}\left(1{\color[rgb]{0,0,0}+\left|{\Upsilon(s,x)}\right|}+\left|{\partial_{x}\varphi(s,x,\tilde{x})}\right|\right)\,\sup_{t\leq s}\left|{x(t)-\tilde{x}(t)}\right|\\ &\qquad\qquad{\color[rgb]{0,0,0}+L_{H}\left(1+\sup\nolimits_{t\leq s}\left|{\tilde{x}(t)}\right|\right)\,\left|{\partial_{x}\varphi(s,x,\tilde{x}){\color[rgb]{0,0,0}+}\partial_{\tilde{x}}\varphi(s,x,\tilde{x})}\right|}\geq 0\end{split}$
whenever $w(s,x,\tilde{x})>0$.
Lemma 5.2.
Let $L\geq 0$, $u$ be a minimax $L$-subsolution and $v$ be a minimax $L$-supersolution of (4.1). Then $(t,x,\tilde{x})\mapsto w(t,x,\tilde{x}):=u(t,x)-v(t,\tilde{x})$, $[0,T]\times\Omega\times\Omega\to{\mathbb{R}}$, is a viscosity $L$-subsolution of (5.2) with parameter $\Upsilon=u$.
Proof.
Let $(s_{0},x_{0},\tilde{x}_{0})\in[0,T)\times\Omega\times\Omega$, $w(s_{0},x_{0},\tilde{x}_{0})>0$, $\varphi\in\underline{\mathcal{A}}^{L}w(s_{0},x_{0},\tilde{x}_{0})$, and $T_{0}\in(s_{0},T]$ such that (5.1) holds. Let $y_{0}:=u(s_{0},x_{0})$, $\tilde{y}_{0}:=v(s_{0},x_{0})$, $z:=\partial_{x}\varphi(s_{0},x_{0},\tilde{x}_{0})$, and $\tilde{z}:=-\partial_{\tilde{x}}\varphi(s_{0},x_{0},\tilde{x}_{0})$. By the minimax semisolution properties of $u$ and $v$, there exist $(x,y)\in\mathcal{Y}^{L}(s_{0},x_{0},y_{0},z)$ and $(\tilde{x},\tilde{y})\in\mathcal{Y}^{L}(s_{0},\tilde{x}_{0},\tilde{y}_{0},\tilde{z})$, such that, for every $t\in[s_{0},T]$,
(5.4) $\begin{split}&[u(t,x)-y_{0}]-[v(t,\tilde{x})-\tilde{y}_{0}]\geq[y(t)-y_{0}]-[\tilde{y}(t)-\tilde{y}_{0}]\\ &\qquad=\int_{s_{0}}^{t}\left[(x^{\prime}(s),z)-H(s,x,y(s),z)-(\tilde{x}^{\prime}(s),\tilde{z})+H(s,\tilde{x},\tilde{y}(s),\tilde{z})\right]\,{\mathrm{d}}s.\end{split}$
By (5.1), $(\varphi-w)(s_{0},x_{0},\tilde{x}_{0})\leq(\varphi-w)(t,x,\tilde{x})$ for every $t\in[s_{0},T_{0}]$. Thus, the chain rule applied to $\varphi$ together with (5.4) yields
$\int_{s_{0}}^{t}\left[\partial_{t}\varphi(s,x,\tilde{x})+(x^{\prime}(s),\partial_{x}\varphi(s,x,\tilde{x}))+(\tilde{x}^{\prime}(s),\partial_{\tilde{x}}\varphi(s,x,\tilde{x}))\right]\,{\mathrm{d}}s$
$\qquad\geq w(t,x,\tilde{x})-w(s_{0},x_{0},\tilde{x}_{0})$
$\qquad\geq\int_{s_{0}}^{t}\left[(x^{\prime}(s),z)-(\tilde{x}^{\prime}(s),\tilde{z})-H(s,x,y(s),z)+H(s,\tilde{x},\tilde{y}(s),\tilde{z})\right]\,{\mathrm{d}}s$
for every $t\in[s_{0},T_{0}]$. Next, let $\delta>0$ be sufficiently small such that $y(s)>\tilde{y}(s)$ for all $s\in[s_{0},s_{0}+\delta]$. This is possible because $w(s_{0},x_{0},\tilde{x}_{0})>0$ and the functions $y$ and $\tilde{y}$ are continuous. Hence, by Assumption 4.2,
(5.5) $\begin{split}0&\leq\int_{s_{0}}^{s_{0}+\delta}\Bigl{[}\partial_{t}\varphi(s,x,\tilde{x})+H(s,x,y(s),z)-H(s,\tilde{x},\tilde{y}(s),\tilde{z})\\ &\quad\quad+(x^{\prime}(s),\partial_{x}\varphi(s,x,\tilde{x})-z)+(\tilde{x}^{\prime}(s),\partial_{\tilde{x}}\varphi(s,x,\tilde{x})+\tilde{z})\Bigr{]}\,{\mathrm{d}}s\\ &\leq\int_{s_{0}}^{s_{0}+\delta}\Bigl{\{}\partial_{t}\varphi(s,x,\tilde{x})+[H(s,x,y(s),z)-H(s,\tilde{x},y(s),z)]\\ &\quad\quad+[H(s,\tilde{x},y(s),z)-H(s,\tilde{x},y(s),\tilde{z})]\\ &\quad\quad+[H(s,\tilde{x},y(s),\tilde{z})-H(s,\tilde{x},\tilde{y}(s),\tilde{z})]\\ &\quad\quad+(x^{\prime}(s),\partial_{x}\varphi(s,x,\tilde{x})-z)+(\tilde{x}^{\prime}(s),\partial_{\tilde{x}}\varphi(s,x,\tilde{x})+\tilde{z})\Bigr{\}}\,{\mathrm{d}}s\\ &\leq\int_{s_{0}}^{s_{0}+\delta}\Bigl{\{}\partial_{t}\varphi(s,x,\tilde{x})+M_{L}(1+\left|{y(s)}\right|+\left|{z}\right|)\sup_{t\leq s}\left|{x(t)-\tilde{x}(t)}\right|\\ &\quad\quad+L_{H}(1+\sup_{t\leq s}\left|{\tilde{x}(t)}\right|)\left|{z-\tilde{z}}\right|+0\\ &\quad\quad+(x^{\prime}(s),\partial_{x}\varphi(s,x,\tilde{x})-z)+(\tilde{x}^{\prime}(s),\partial_{\tilde{x}}\varphi(s,x,\tilde{x})+\tilde{z})\Bigr{\}}\,{\mathrm{d}}s.\end{split}$
Finally, dividing (5.5) by $\delta$ and letting $\delta\downarrow 0$ yields (5.3). ∎
Theorem 5.3.
Fix $L\geq 0$. Let $w:[0,T]\times\Omega\times\Omega\to{\mathbb{R}}$ be a viscosity $L$-subsolution of (5.2) with an upper semi-continuous parameter $\Upsilon$. Suppose that $w(T,x,x)\leq 0$ for every $x\in\Omega$. Then $w(t,x,x)\leq 0$ for every $(t,x)\in[0,T]\times\Omega$.
Proof.
Assume that there is a point $(s_{0},x_{0})\in[0,T)\times\Omega$ such that
$M_{0}:=w(s_{0},x_{0},x_{0})>0.$
Proceeding along the lines of the proof of Theorem 4.2 in [BK18JFA], we will obtain a contradiction. The main difference compared to [BK18JFA] is the choice of a different penalty functional. Here, we use $\Psi:[s_{0},T]\times\Omega\times\Omega\to{\mathbb{R}}$ defined by
$\Psi(s,x,\tilde{x}):=\begin{cases}\frac{\sup\limits_{t\leq s}\left|{x(t)-\tilde{x}(t)}\right|^{2}-\left|{x(s)-\tilde{x}(s)}\right|^{2}}{\sup\limits_{t\leq s}\left|{x(t)-\tilde{x}(t)}\right|^{2}}+\left|{x(s)-\tilde{x}(s)}\right|^{2}&\text{if $\sup\limits_{t\leq s}\left|{x(t)-\tilde{x}(t)}\right|>0$,}\\ 0&\text{otherwise.}\end{cases}$
This functional has been introduced in [zhou2020viscosity] (see also [GLP21AMO] for further details). By the arguments in [GLP21AMO, Appendix B], one can deduce that $\Psi\in C^{1,1,1}([s_{0},T]\times\Omega\times\Omega)$ with derivatives $\partial_{t}\Psi(s,x,\tilde{x})=0$ and
$\partial_{x}\Psi(s,x,\tilde{x})=-\partial_{\tilde{x}}\Psi(s,x,\tilde{x})$
=$=\begin{cases}\left(2-\frac{4\sup\limits_{t\leq s}\left|{x(t)-\tilde{x}(t)}\right|^{2}-\left|{x(s)-\tilde{x}(s)}\right|^{2}}{\sup\limits_{t\leq s}\left|{x(t)-\tilde{x}(t)}\right|^{2}}\right)\cdot[x(s)-\tilde{x}(s)]&\text{if $\sup\limits_{t\leq s}\left|{x(t)-\tilde{x}(t)}\right|>0$,}\\ 0&\text{otherwise.}\end{cases}$
Given ${\varepsilon}>0$, define $\Phi_{\varepsilon}:[s_{0},T]\times\Omega\times\Omega\to{\mathbb{R}}$ by
$\Phi_{\varepsilon}(s,x,\tilde{x}):=w(s,x,\tilde{x})-\frac{T-s}{T-s_{0}}\cdot\frac{M_{0}}{2}-\frac{1}{{\varepsilon}}\Psi(s,x,\tilde{x}).$
Fix a point $k_{\varepsilon}=(s_{\varepsilon},x_{\varepsilon},\tilde{x}_{\varepsilon})$ at which the u.s.c. map $\Phi_{\varepsilon}$ attains a maximum on the compact set $K:=[s_{0},T]\times\mathcal{X}^{L}(s_{0},x_{0})\times\mathcal{X}^{L}(s_{0},x_{0})$. Note that
$M_{\varepsilon}:=\Phi_{\varepsilon}(k_{\varepsilon})\geq\Phi_{\varepsilon}(s_{0},x_{0},x_{0})=w(s_{0},x_{0},x_{0})-\frac{M_{0}}{2}=\frac{M_{0}}{2}.$
Thus
(5.6) $w(k_{\varepsilon})\geq\frac{M_{0}}{2}+\frac{T-s_{\varepsilon}}{T-s_{0}}\cdot\frac{M_{0}}{2}+\frac{1}{{\varepsilon}}\Psi(k_{\varepsilon})\geq\frac{M_{0}}{2}>0.$
Moreover, $s_{\varepsilon}<T$ if ${\varepsilon}$ is sufficiently small (cf. Proposition 3.7 and its proof in [CrandallIshiiLions]).
Define a map $\varphi_{\varepsilon}:[s_{\varepsilon},T]\times\Omega\times\Omega\to{\mathbb{R}}$ by
$\varphi_{\varepsilon}(s,x,\tilde{x}):=M_{\varepsilon}+\frac{T-s}{T-s_{0}}\cdot\frac{M_{0}}{2}+\frac{1}{{\varepsilon}}\Psi(s,x,\tilde{x}).$
Then $\varphi_{\varepsilon}\in\underline{\mathcal{A}}^{L}w(k_{\varepsilon})$ with corresponding time $T_{0}=T$ because
$\varphi_{\varepsilon}(k_{\varepsilon})-w(k_{\varepsilon})=\overbrace{w(k_{\varepsilon})-\frac{T-s_{\varepsilon}}{T-s_{0}}\cdot\frac{M_{0}}{2}-\frac{1}{{\varepsilon}}\Psi(k_{\varepsilon})}^{=M_{\varepsilon}}$
$\qquad\qquad\qquad\qquad\qquad\qquad+\frac{T-s_{\varepsilon}}{T-s_{0}}\cdot\frac{M_{0}}{2}+\frac{1}{{\varepsilon}}\Psi(k_{\varepsilon})-w(k_{\varepsilon})=0$
$\qquad\leq M_{\varepsilon}-\Phi_{\varepsilon}(s,x,\tilde{x})=M_{\varepsilon}+\frac{T-s}{T-s_{0}}\cdot\frac{M_{0}}{2}+\frac{1}{{\varepsilon}}\Psi(s,x,\tilde{x})-w(s,x,\tilde{x})$
$\qquad=\varphi_{\varepsilon}(s,x,\tilde{x})-w(s,x,\tilde{x})$
for every $(s,x,\tilde{x})\in[s_{\varepsilon},T]\times\mathcal{X}^{L}(s_{\varepsilon},x_{\varepsilon})\times\mathcal{X}^{L}(s_{\varepsilon},\tilde{x}_{\varepsilon})$. Note that
$\partial_{t}\varphi_{\varepsilon}(s,x,\tilde{x})=-\frac{M_{0}}{2(T-s_{0})}\quad\text{and}\quad\partial_{x}\varphi_{\varepsilon}(s,x,\tilde{x})=-\partial_{\tilde{x}}\varphi_{\varepsilon}(s,x,\tilde{x})=\frac{1}{{\varepsilon}}\partial_{x}\Psi(s,x,\tilde{x}).$
Consequently, since $w$ is a viscosity $L$-subsolution of (5.2) and since (5.6) holds,
(5.7) $\begin{split}0&\leq-\frac{M_{0}}{2(T-s_{0})}+M_{L}\left(1+\left|{\Upsilon(s_{\varepsilon},x_{\varepsilon})}\right|+\frac{1}{{\varepsilon}}\left|{\partial_{x}\Psi(k_{\varepsilon})}\right|\right)\cdot\sup_{t\leq s_{\varepsilon}}\left|{x_{\varepsilon}(t)-\tilde{x}_{\varepsilon}(t)}\right|\\ &\leq-\frac{M_{0}}{2(T-s_{0})}+\tilde{M}_{L}\left(1+\frac{1}{{\varepsilon}}\left|{\partial_{x}\Psi(k_{\varepsilon})}\right|\right)\cdot\sup_{t\leq s_{\varepsilon}}\left|{x_{\varepsilon}(t)-\tilde{x}_{\varepsilon}(t)}\right|\end{split}$
for some constant $\tilde{M}_{L}>0$ independent from ${\varepsilon}$. The second line of (5.7) follows from the upper semi-continuity of $\Upsilon$ and the compactness of $\mathcal{X}^{L}(s_{0},x_{0})$. It remains to note that (cf. [GLP21AMO, page S1102])
$\left|{\partial_{x}\Psi(s,x,\tilde{x})}\right|\leq 2\left|{x(s)-\tilde{x}(s)}\right|$
and also, by [zhou2020viscosity, Lemma 2.3],
$\frac{3-\sqrt{5}}{2}\,\sup_{t\leq s}\left|{x(t)-\tilde{x}(t)}\right|^{2}\leq\Psi(s,x,\tilde{x})\leq 2\,\sup_{t\leq s}\left|{x(t)-\tilde{x}(t)}\right|^{2}$
for every $(s,x,\tilde{x})\in[0,T]\times\Omega\times\Omega$. Then we can deduce, as in the proof of Theorem 4.2 in [BK18JFA], that
$\left(1+\frac{1}{{\varepsilon}}\left|{\partial_{x}\Psi(k_{\varepsilon})}\right|\right)\cdot\sup_{t\leq s_{\varepsilon}}\left|{x_{\varepsilon}(t)-\tilde{x}_{\varepsilon}(t)}\right|\leq\sqrt{\Psi(k_{\varepsilon})}+\frac{C}{{\varepsilon}}\Psi(k_{\varepsilon})\to 0\quad\text{as ${\varepsilon}\downarrow 0$}$
for some constant $C>0$, which contradicts (5.7). ∎
Lemma 5.2 and Theorem 5.3 immediately yield the following result.
Corollary 5.4.
Let $L\geq 0$, $u$ be a minimax $L$-subsolution and $v$ be a minimax $L$-supersolution of (4.1). Then $u\leq v$ on $[0,T]\times\Omega$.
6. Existence
We show existence of minimax solutions via Perron’s method (see [BardiCapuzzoDolcetta, section V.2] regarding a corresponding treatment for non-continuous viscosity solutions in the non-path-dependent case). More precisely, the scheme888 This scheme, which provides existence for minimax solutions, actually predates “Perron’s method” for viscosity solutions introduced in [Ishii87Perron] (we refer to the discussion in [Subbotin, section 10.8] for more details). in [Subbotin, section 8] is adapted to path-dependent Hamilton-Jacobi equations. In the case of Hamiltonians without $u$-dependence, this has already been done in [Lukoyanov03, section 7] (cf. also [BK18JFA, section 5], whose structure we follow here).
Definition 6.1.
Let $L\geq 0$ and $u:[0,T]\times\Omega\to[-\infty,\infty]$.
(i) $u$ is a non-continuous minimax $L$-supersolution of (4.1) if $u(T,\cdot)=h$ on $\Omega$ and, for every $(t_{0},x_{0},z)\in[0,T)\times\Omega\times{\mathbb{R}}^{d}$, $y_{0}>u(t_{0},x_{0})$, and $T_{0}\in(t_{0},T]$,
(6.1) $\inf_{(x,y)\in\mathcal{Y}^{L}(t_{0},x_{0},y_{0},z)}\{u(T_{0},x)-y(T_{0})\}\leq 0.$
ii) $u$ is a non-continuous minimax $L$-subsolution of (4.1) if $u(T,\cdot)=h$ on $\Omega$ and, for every $(t_{0},x_{0},z)\in[0,T)\times\Omega\times{\mathbb{R}}^{d}$, $y_{0}<u(t_{0},x_{0})$, and $T_{0}\in(t_{0},T]$,
(6.2) $\sup_{(x,y)\in\mathcal{Y}^{L}(t_{0},x_{0},y_{0},z)}\{u(T_{0},x)-y(T_{0})\}\geq 0.$
(iii) $u$ is a non-continuous minimax $L$-solution of (4.1) if it is both, a non-continuous minimax $L$-super- and a non-continuous minimax $L$-subsolution of (4.1)
First, we establish existence of non-continuous minimax supersolutions. Following [Subbotin, Proposition 8.6], we define functions $\mu^{z}_{+}:[0,T]\times\Omega\times{\mathbb{R}}\to[-\infty,\infty]$ and $u^{z}_{+}:[0,T]\times\Omega\to[-\infty,\infty]$, $z\in{\mathbb{R}}^{d}$, by
$\mu^{z}_{+}(t_{0},x_{0},y_{0})$ $:=\sup_{(x,y)\in\mathcal{Y}^{(L_{H})}(t_{0},x_{0},y_{0},z)}\{h(x)-y(T)\},$
$u^{z}_{+}(t_{0},x_{0})$ $:=\sup\{r\in{\mathbb{R}}:\,\mu^{z}_{+}(t_{0},x_{0},r)\geq 0\}.$
Lemma 6.2.
Let $z\in{\mathbb{R}}^{d}$. Then $u^{z}_{+}$ is a non-continuous minimax $L_{H}$-supersolution of (4.1) and it is [,)[$\infty$,$\infty$)[ - ∞ , ∞ )-valued.
Proof.
(i) Boundary condition: Let $t_{0}=T$ and $x_{0}\in\Omega$. Since $\mu^{z}_{+}(T,x_{0},y_{0})=h(x_{0})-y_{0}$ for all $y_{0}\in{\mathbb{R}}$, we have $u^{z}_{+}(T,x_{0})=\sup\{r\in{\mathbb{R}}:\,h(x_{0})-r\geq 0\}=h(x_{0})$.
(ii) Interior condition: Let $(t_{0},x_{0},\tilde{z})\in[0,T)\times\Omega\times{\mathbb{R}}^{d}$, $y_{0}>u^{z}_{+}(t_{0},x_{0})$, and $T_{0}\in(t_{0},T]$. Also pick a pair $(\tilde{x},\tilde{y})\in\mathcal{Y}^{L_{H}}(t_{0},x_{0},y_{0},z)\cap\mathcal{Y}^{L_{H}}(t_{0},x_{0},y_{0},\tilde{z})$. This is possible according to Remark 4.3. To verify the interior minimax supersolution property, it suffices to show that $u^{z}_{+}(T_{0},\tilde{x})\leq\tilde{y}(T_{0})$. To this end, note first that $\mu^{z}_{+}(t_{0},x_{0},y_{0})<0$ because otherwise $u^{z}_{+}(t_{0},x_{0})\geq y_{0}$, which would contradict $y_{0}>u^{z}_{+}(t_{0},x_{0})$. Therefore
(6.3) $\begin{split}0&>\mu^{z}_{+}(t_{0},x_{0},y_{0})\\ &\geq\sup\{h(x)-y(T):\,(x,y)\in\mathcal{Y}^{(L_{H})}(t_{0},x_{0},y_{0},z)\\ &\qquad\qquad\qquad\qquad\qquad\qquad\text{ and $(x,y)|_{[0,T_{0}]}=(\tilde{x},\tilde{y})|_{[0,T_{0}]}$}\}\\ &=\sup\{h(x)-y(T):\,(x,y)\in\mathcal{Y}^{(L_{H})}(T_{0},\tilde{x},\tilde{y}(T_{0}),z)\}\\ &=\mu^{z}_{+}(T_{0},\tilde{x},\tilde{y}(T_{0})).\end{split}$
Let us write $u^{z}_{+}(T_{0},\tilde{x})=\sup R$, where $R:=\{r\in{\mathbb{R}}:\,\mu^{z}_{+}(T_{0},\tilde{x},r)\geq 0\}$. By (6.3), $\tilde{y}(T_{0})\not\in R$. Note that, for every $s\geq 0$, we have
(6.4) $\mu^{z}_{+}(T_{0},\tilde{x},\tilde{y}(T_{0})+s)\leq\mu^{z}_{+}(T_{0},\tilde{x},\tilde{y}(T_{0}))-s\leq 0$
(this follows from Assumption 4.2 (v) and can be shown exactly as equation (8.4) in [Subbotin]). Thus $R\subseteq[-\infty,\tilde{y}(T_{0}))$, i.e., $u^{z}_{+}(T_{0},\tilde{x})\leq\tilde{y}(T_{0})$.
(iii) $u^{z}_{+}$ is [,)[$\infty$,$\infty$)[ - ∞ , ∞ )-valued: First, note that $\mu^{z}_{+}$ is [,)[$\infty$,$\infty$)[ - ∞ , ∞ )-valued due to the compactness of the sets $\mathcal{Y}^{L}(t_{0},x_{0},y_{0},z)$ (Remark 4.3) and the continuity of $h$ (Assumption 4.1). Now assume that $u^{z}_{+}(t_{0},x_{0})=\infty$ for some $(t_{0},x_{0})\in[0,T]\times\Omega$. Then there is an increasing sequence $(r_{n})_{n}$ in ${\mathbb{R}}$ with $\mu^{z}_{+}(t_{0},x_{0},r_{n})\geq 0$ and $r_{n}\to\infty$ as $n\to\infty$. But, by (6.4),
$0\leq\mu^{z}_{+}(t_{0},x_{0},r_{n})=\mu^{z}_{+}(t_{0},x_{0},r_{1}+(r_{n}-r_{1}))\leq\mu^{z}_{+}(t_{0},x_{0},r_{1})-(r_{n}-r_{1})\to-\infty$
as $n\to\infty$. This contradicts our assumption and thus concludes the proof. ∎
Define $u_{0}:[0,T]\times\Omega\to[-\infty,\infty]$ by
[…]
Proposition 6.3.
The function $u_{0}$ is a non-continuous minimax $L_{H}$-solution of (4.1) and it is ${\mathbb{R}}$-valued.
Theorem 6.4.
The function $u_{0}$ is the unique minimax $L_{H}$-solution of (4.1).
6.1. Proof of Proposition 6.3
The proof consists of four parts.
(i) $u_{0}$ is ${\mathbb{R}}$-valued: By the definition of $u_{0}$ and Lemma 6.2, $u_{0}$ is [,)[$\infty$,$\infty$)[ - ∞ , ∞ )-valued. To show that $u_{0}$ is (,]($\infty$,$\infty$]( - ∞ , ∞ ]-valued, consider, following [Subbotin, Proposition 8.4] functions $\mu^{z}_{-}:[0,T]\times\Omega\times{\mathbb{R}}\to[-\infty,\infty]$ and $u^{z}:[0,T]\times\Omega\to[-\infty,\infty]$, $z\in{\mathbb{R}}^{d}$, defined by
$\mu^{z}_{-}(t_{0},x_{0},y_{0})$ $:=\inf_{(x,y)\in\mathcal{Y}^{(L_{H})}(t_{0},x_{0},y_{0},z)}\{h(x)-y(T)\},$
$u^{z}_{-}(t_{0},x_{0})$ $:=\inf\{r\in{\mathbb{R}}:\,\mu^{z}_{-}(t_{0},x_{0},r)\leq 0\}.$
Fix $z\in{\mathbb{R}}^{d}$. Note that one can show similarly as (6.4) that, for all $s\geq 0$, we have
$\mu^{z}_{-}(t_{0},x_{0},y_{0}+s)\leq\mu^{z}_{-}(t_{0},x_{0},y_{0})-s,$
i.e., $u^{z}_{-}$ is (,]($\infty$,$\infty$]( - ∞ , ∞ ]-valued (cf. part (iii) of the proof of Lemma 6.2). Next, let $u$ be an arbitrary non-continuous minimax $L_{H}$-solution of (4.1). We show that, for every $(t,x)\in[0,T]\times\Omega$, we have $u(t,x)\geq u^{z}_{-}(t,x)$, which then immediately yields $u_{0}(t,x)\geq u^{z}_{-}(t,x)>-\infty$. To see this, assume that $u(t_{0},x_{0})<u^{z}_{-}(t_{0},x_{0})$ for some $(t_{0},x_{0})\in[0,T]\times\Omega$. Then we can pick a $y_{0}\in(u(t_{0},x_{0}),u^{z}_{-}(t_{0},x_{0}))$ and thus, by (6.1) with $T_{0}=T$ and $u(T,\cdot)=h$, we have $\mu^{z}_{-}(t_{0},x_{0},y_{0})\leq 0$. But this implies $u^{z}_{-}(t_{0},x_{0})\leq y_{0}<u^{z}_{-}(t_{0},x_{0})$. We can conclude that $u_{0}$ is ${\mathbb{R}}$-valued.
(ii) Boundary condition: By the definition of $u_{0}$ and Definition 6.1, $u_{0}(T,\cdot)=h$.
(iii) Interior minimax supersolution property: Let $(t_{0},x_{0},z)\in[0,T)\times\Omega\times{\mathbb{R}}^{d}$, $y_{0}>u_{0}(t_{0},x_{0})$, and $T_{0}\in(t_{0},T]$. By the definition of $u_{0}$ and Lemma 6.2, there exists a non-continuous minimax $L_{H}$-supersolution $u$ of (4.1) such that $y_{0}>u(t_{0},x_{0})$. Thus, for every $n\in{\mathbb{N}}$, there is a pair $(x_{n},y_{n})\in\mathcal{Y}^{L_{H}}(t_{0},x_{0},y_{0},z)$ such that $u(T_{0},x_{n})\leq y_{n}(T_{0})+n^{-1}$. Hence $u_{0}(T_{0},x_{n})\leq y_{n}(T_{0})+n^{-1}$, i.e., $u_{0}$ is a non-continuous minimax $L_{H}$-supersolution of (4.1).
(iv) Interior minimax subsolution property: Let $(t_{1},x_{1},z)\in[0,T)\times\Omega\times{\mathbb{R}}^{d}$, $y_{1}<u_{0}(t_{1},x_{1})$, and $T_{1}\in(t_{1},T]$. We need to show that
(6.5) $\sup_{(x,y)\in\mathcal{Y}^{L_{H}}(t_{1},x_{1},y_{1},z)}\{u_{0}(T_{1},x)-y(T_{1})\}\geq 0.$
To this end, we consider, following [Subbotin, Proposition 8.5], the functions $\mu^{z}:[0,T_{1}]\times\Omega\times{\mathbb{R}}\to[-\infty,\infty]$ and $u^{z}:[0,T]\times\Omega\to[-\infty,\infty]$ defined by
$\mu^{z}(t_{0},x_{0},y_{0})$ $:=\sup_{(x,y)\in\mathcal{Y}^{L_{H}}(t_{0},x_{0},y_{0},z)}\{u_{0}(T_{1},x)-y(T_{1})\},$
$u^{z}(t_{0},x_{0})$ :=$:=\begin{cases}\sup\{r\in{\mathbb{R}}:\,\mu^{z}(t_{0},x_{0},r)\geq 0\}\quad\text{if $t_{0}\leq T_{1}$,}\\ u_{0}(t_{0},x_{0})\qquad\qquad\qquad\qquad\quad\,\,\,\,\text{if $t_{0}>T_{1}$.}\end{cases}$
Assume momentarily that
(6.6) $u^{z}$ is a non-continuous minimax $L_{H}$-supersolution of (4.1).
Then, by the definition of $u_{0}$, by part (i) of this proof, and by noting that $t_{0}<T_{1}$, we have $-\infty<u_{0}(t_{1},x_{1})\leq u^{z}(t_{1},x_{1})=\sup\{r\in{\mathbb{R}}:\mu^{z}(t_{1},x_{1},r)\geq 0\}$. Thus, recalling that $y_{1}<u_{0}(t_{1},x_{1})$, we can see that there exists an $r_{1}\in(y_{0},\infty)$ such that $\mu^{z}(t_{1},x_{1},r_{1})\geq 0$ and then after showing similarly as (6.4) that
$\mu^{z}(t_{1},x_{1},y_{1}+(r_{1}-y_{1}))\leq\mu^{z}(t_{1},x_{1},y_{1})-(r_{1}-y_{0})$
holds, we obtain (6.5). Hence, it remains to establish (6.6), which we do next.
(iv) (a) Boundary condition for $u^{z}$:
Case 1. Let $T_{1}=T$. Then $\mu^{z}(T,x_{0},y_{0})=h(x_{0})-y_{0}$ and thus $u^{z}(T,x_{0})=h(x_{0})$.
Case 2. Let $T_{1}<T$. Then $u^{z}(T,x_{0})=u_{0}(T,x_{0})=h(x_{0})$.
(iv) (b) Interior condition for $u^{z}$: Fix $(t_{0},x_{0})\in[0,T)\times\Omega$, $y_{0}>u^{z}(t_{0},x_{0})$, $\tilde{z}\in{\mathbb{R}}^{d}$, and $T_{0}\in(t_{0},T]$. We need to show
(6.7) $\inf_{(x,y)\in\mathcal{Y}^{L_{H}}(t_{0},x_{0},y_{0},\tilde{z})}\{u^{z}(T_{0},x)-y(T_{0})\}\leq 0.$
Case 1. Let $T_{1}<t_{0}$. Then $u_{0}=u^{z}$ on $[t_{0},T]$ and (6.7) follows from $u_{0}$ being a non-continuous minimax $L_{H}$-supersolution of (4.1) (see part (iii) of this proof).
Case 2. Let $T_{1}=t_{0}$. Then $\mu^{z}(t_{0},x_{0},r)=u_{0}(t_{0},x_{0})-r$. Thus $u_{0}=u^{z}$ on $[t_{0},T]$ and (6.7) follows as in Case 1 of part (iv) (b) of this proof.
Case 3. Let $t_{0}<T_{0}\leq T_{1}$. To obtain (6.7), follow the proof of part (ii) of Lemma 6.2 but replace $h$ by $u_{0}(T_{1},\cdot)$ and $T$ by $T_{1}$.
Case 4. Let $t_{0}<T_{1}\leq T_{0}$. Then $u^{z}(T_{0},\cdot)=u_{0}(T_{0},\cdot)$ and thus (6.7) follows by noting that
$\inf_{(x,y)\in\mathcal{Y}^{L_{H}}(t_{0},x_{0},y_{0},\tilde{z})}\{u_{0}(T_{0},x)-y(T_{0})\}\leq\inf_{(x,y)\in\mathcal{Y}^{L_{H}}(T_{1},\tilde{x},\tilde{y}(T_{1}),\tilde{z})}\{u_{0}(T_{0},x)-y(T_{0})\},$
where $(\tilde{x},\tilde{y})\in\mathcal{Y}^{L_{H}}(t_{0},x_{0},z)\cap\mathcal{Y}^{L_{H}}(t_{0},x_{0},\tilde{z})$ (cf. Remark 4.3), and that $u_{0}$ is a a non-continuous minimax $L_{H}$-supersolution of (4.1) (see part (iii) of this proof) together with $\tilde{y}(T_{1})>u_{0}(T_{1},\tilde{x})$. Note that if the last inequality was not true, then $\mu^{z}(t_{0},x_{0},y_{0})\geq 0$ and thus $y_{0}\leq u^{z}(t_{0},x_{0})$, which would be a contradiction.
This concludes the proof of Proposition 6.3. ∎
6.2. Proof of Theorem 6.4
Consider the u.s.c. envelope $(u_{0})^{\ast}:[0,T]\times\Omega\to[-\infty,\infty]$ and the l.s.c. envelope $(u_{0})_{\ast}:[0,T]\times\Omega\to[-\infty,\infty]$, which are defined by
$(u_{0})^{\ast}(t_{0},x_{0})$ $:=\inf_{\delta>0}\sup_{(t,x)\in O_{\delta}(t_{0},x_{0})}u_{0}(t,x),$
$(u_{0})_{\ast}(t_{0},x_{0})$ $:=\sup_{\delta>0}\inf_{(t,x)\in O_{\delta}(t_{0},x_{0})}u_{0}(t,x),$
where $O_{\delta}(t_{0},x_{0})=\{(t,x)\in[0,T]\times\Omega:\,\mathbf{d}_{\infty}((t_{0},x_{0}),(t_{n},x_{n}))<\delta\}$. We show that $(u_{0})_{\ast}$ is a minimax $L_{H}$-supersolution of (4.1) and $(u_{0})^{\ast}$ is a minimax $L_{H}$-subsolution of (4.1). The comparison principle (Corollary 5.4) yields $(u_{0})^{\ast}\leq(u_{0})_{\ast}$. But, since, by definition, $(u_{0})_{\ast}\leq u_{0}\leq(u_{0})^{\ast}$ holds, we can conclude that $u_{0}$ is a minimax $L_{H}$-subsolution of (4.1). Again, by Corollary 5.4, it is the only one.
Thus it remains to establish the minimax semisolution properties of the semicontinuous envelopes of $u_{0}$.
(i) $(u_{0})^{\ast}$ is (,]($\infty$,$\infty$]( - ∞ , ∞ ]- and $(u_{0})_{\ast}$ is [,)[$\infty$,$\infty$)[ - ∞ , ∞ )-valued: This follows from $u_{0}$ being ${\mathbb{R}}$-valued (Proposition 6.3) and $(u_{0})_{\ast}\leq u_{0}\leq(u_{0})^{\ast}$.
(ii) $(u_{0})^{\ast}$ is [,)[$\infty$,$\infty$)[ - ∞ , ∞ )- and $(u_{0})_{\ast}$ is (,]($\infty$,$\infty$]( - ∞ , ∞ ]-valued: We only show that $(u_{0})^{\ast}<\infty$. The remaining part can be shown similarly. Let $(t_{0},x_{0})\in[0,T]\times\Omega$ and $(t_{n},x_{n})_{n}$ be a sequence that converges to $(t_{0},x_{0})$ with respect to $\mathbf{d}_{\infty}$ such that $(u_{0})^{\ast}(t_{0},x_{0})=\lim_{n}u_{0}(t_{n},x_{n})$. Fix $z\in{\mathbb{R}}^{d}$. By the definition of $u_{0}$ and Lemma 6.2, we have $\lim_{n}u_{0}(t_{n},x_{n})\leq\lim_{n}u^{z}_{+}(t_{n},x_{n})=\lim_{n}r_{n}$ for some sequence $(r_{n})_{n}$ in ${\mathbb{R}}$ with $\mu^{z}_{+}(t_{n},x_{n},r_{n})\geq 0$. By Remark 4.3 and continuity of $h$, for each $n\in{\mathbb{N}}$, there is a pair $(\tilde{x}_{n},\tilde{y}_{n})\in\mathcal{Y}^{L_{H}}(t_{n},x_{n},r_{n},z)$ such that $\mu^{z}_{+}(t_{n},x_{n},r_{n})=h(\tilde{x}_{n})-\tilde{y}_{n}(T)$. Without loss of generality, we can, by Remark 2.2, assume that $(\tilde{x}_{n})_{n}$ converges to some $\tilde{x}_{0}\in\mathcal{X}^{L_{H}}(t_{0},x_{0})$. Assume now for the sake of a contradiction that $(u_{0})^{\ast}(t_{0},x_{0})=\infty$. Then we can assume, without loss of generality, that $(r_{n})_{n}$ is strictly increasing and strictly positive. This in turn implies the following. Given $n\in{\mathbb{N}}$, fix a solution $y^{0}_{n}\in C([t_{n},T])$ of $y^{0}_{n}(t)=\int_{t_{n}}^{t}[(\tilde{x}_{n}^{\prime}(s),z)-H(s,\tilde{x}_{n},y^{0}_{n}(s),z)]\,{\mathrm{d}}s$, $t\in[t_{n},T]$. Then, we have $\tilde{y}_{n}>y^{0}_{n}$ on $(t_{n},T)$, as otherwise, by continuity of $y^{0}_{n}$ and $\tilde{y}_{n}$, there is a smallest time $\tau_{n}\in(t_{n},T]$ such that $\tilde{y}_{n}(\tau)=y^{0}_{n}(\tau_{n})$ and $\tilde{y}_{n}>y^{0}_{n}$ on $(t_{n},\tau_{n})$ but then
00 0 $=\tilde{y}_{n}(\tau)-y^{0}_{n}(\tau_{n})=r_{n}+\int^{\tau}_{t_{n}}\left[H(s,\tilde{x}_{n},y^{0}_{n}(s),z)-H(s,\tilde{x}_{n},\tilde{y}_{n}(s),z)\right]\,{\mathrm{d}}s\geq r_{n}>0$
(cf. [Subbotin, page 72]), which is absurd. Using now $\tilde{y}_{n}>y^{0}_{n}$ together with $h(\tilde{x}_{n})-\tilde{y}_{n}(T)\geq 0$, we obtain
(6.8) $\begin{split}r_{n}&\leq h(\tilde{x}_{n})+\int_{t_{n}}^{T}\left[H(s,\tilde{x}_{n},\tilde{y}_{n}(s),z)-(\tilde{x}^{\prime}_{n}(s),z)\right]\,{\mathrm{d}}s\\ &\leq h(\tilde{x}_{n})+\int_{t_{n}}^{T}\left[H(s,\tilde{x}_{n},y^{0}_{n}(s),z)-(\tilde{x}^{\prime}_{n}(s),z)\right]\,{\mathrm{d}}s.\end{split}$
Note that, since $(t_{n},x_{n},y^{0}_{n}(t_{n}))=(t_{n},x_{n},0)\to(t_{0},x_{0},0)$ as $n\to\infty$, we can, thanks to Remark 4.4, assume that $(\tilde{x}_{n},{\mathbf{1}}_{[t_{n},T]}\,y_{n}^{0})\to(\tilde{x}_{0},y_{0}^{0})$ as $n\to\infty$ for some $y_{0}^{0}\in D([0,T])$ with $(\tilde{x}_{0},y_{0}^{0}|_{[t_{0},T]})\in\mathcal{Y}^{L}(t_{0},x_{0},0,z)$. Thus, letting $n\to\infty$ in (6.8) yields
$\infty\leq h(\tilde{x}_{0})+\int_{t_{0}}^{T}\left[H(s,\tilde{x}_{0},y^{0}_{0}(s),z)-(\tilde{x}^{\prime}_{0}(s),z)\right]\,{\mathrm{d}}s<\infty,$
which is absurd.999 Alternatively, one can estimate $H$ by using $z=0$ and Assumption 4.2 (vi) to obtain $\infty\leq C<\infty$ for some constant $C$. Therefore, $(u_{0})^{\ast}<\infty$.
(iii) Boundary conditions: We only show $(u_{0})^{\ast}(T,\cdot)\leq h$. Proving $(u_{0})_{\ast}\geq h$ can be done similarly. Let $x_{0}\in\Omega$ and $(t_{n},x_{n})_{n}$ be a sequence in $[0,T]\times\Omega$ that converges to $(T,x_{0})$ and that satisfies $\lim_{n}u_{0}(t_{n},x_{n})=(u_{0})^{\ast}(T,x_{0})$. For each $n\in{\mathbb{N}}$, there is, by Proposition 6.3, an $(\tilde{x}_{n},\tilde{y}_{n})\in\mathcal{Y}^{L_{H}}(t_{n},x_{n},u_{0}(t_{n},x_{n})-1/n,z)$ such that $\tilde{y}_{n}(T)\leq u_{0}(T,\tilde{x}_{n})=h(\tilde{x}_{n})$. By Remark 4.4, we can assume that the sequence $(\tilde{x}_{n},t\mapsto{\mathbf{1}}_{[t_{n},T]}(t)\,\tilde{y}_{n}(t))_{n}$ converges in $(\Omega\times D([0,T]),\lVert{\cdot}\rVert_{\infty})$ to a pair $(\tilde{x},\tilde{y})$ with $(\tilde{x},\tilde{y}|_{\{T\}})\in\mathcal{Y}^{L_{H}}(T,x_{0},(u_{0})^{\ast}(T,x_{0}),z)$, i.e., $\tilde{y}(T)=(u_{0})^{\ast}(T,x_{0})$ and $\tilde{x}=x_{0}$. Thus $(u_{0})^{\ast}(T,x_{0})=\lim_{n}\tilde{y}_{n}(T)\leq\lim_{n}h(\tilde{x}_{n})=h(x_{0})$ thanks to the continuity of $h$ (Assumption 4.1).
(iv) Interior conditions: We only establish the minimax $L_{H}$-subsolution property of $(u_{0})^{\ast}$. Our argument is very close to [Subbotin, page 78]. Let $(t_{0},x_{0},z)\in[0,T)\times\Omega\times{\mathbb{R}}^{d}$, $y_{0}\leq(u_{0})^{\ast}(t_{0},x_{0})$, and $T_{0}\in(t_{0},T]$. Let $(t_{n},x_{n})_{n}$ be a sequence in $[0,T)\times\Omega$ that converges to $(t_{0},x_{0})$ and satisfies $\lim_{n}u_{0}(t_{n},x_{n})=(u_{0})^{\ast}(t_{0},x_{0})$ and $t_{n}<T_{0}$ for every $n\in{\mathbb{N}}$. Next, put $y_{n}:=u_{0}(t_{n},x_{n})-1/n+y_{0}-(u_{0})^{\ast}(t_{0},x_{0})$, $n\in{\mathbb{N}}$, so that $y_{n}<u_{0}(t_{n},x_{n})$ and $(t_{n},x_{n},y_{n})\to(t_{0},x_{0},y_{0})$ as $n\to\infty$. By Proposition 6.3, $u_{0}$ is a non-continuous minimax $L_{H}$-subsolution of (4.1) and thus we obtain from $y_{n}<u_{0}(t_{n},x_{n})$, $n\in{\mathbb{N}}$, the existence of a pairs $(\tilde{x}_{n},\tilde{y}_{n})\in\mathcal{Y}^{L}(t_{n},x_{n},y_{n},z)$ that satisfy $\tilde{y}_{n}(T_{0})-1/n\leq u_{0}(T_{0},\tilde{x}_{n})$. By Remark 4.4 and $\lim_{n}(t_{n},x_{n},y_{n})=(t_{0},x_{0},y_{0})$, we can assume that $(\tilde{x}_{n},t\mapsto{\mathbf{1}}_{[t_{n},T]}(t)\,\tilde{y}_{n}(t))_{n}$ converges in $(\Omega\times D([0,T]),\lVert{\cdot}\rVert_{\infty})$ to some pair $(\tilde{x},\tilde{y})$ with $(\tilde{x},\tilde{y}|_{[t_{0},T]})\in\mathcal{Y}^{L_{H}}(t_{0},x_{0},y_{0},z)$. Thus
$\tilde{y}(T_{0})=\lim_{n}\tilde{y}_{n}(T_{0})\leq\limsup_{n}u_{0}(T_{0},\tilde{x}_{n})\leq(u_{0})^{\ast}(T_{0},\tilde{x}).$
By Lemma 4.7 (or, more precisely, its counterpart for minimax subsolutions), the minimax $L_{H}$-subsolution property of $(u_{0})^{\ast}$ has been established.
This concludes the proof of Theorem 6.4. ∎
7. Optimal control with discount factors
Let $A$ be a topological space. Suppose that $A$ is the countable union of compact metrizable subsets of $A$ (cf. Chapter 21 of [CohenElliott]). Let $\mathcal{A}$ be the set of all Borel measurable functions from $[0,T]$ to $A$. Further data for our control problem are functions $f:[0,T]\times\Omega\times A\to{\mathbb{R}}^{d}$, $\lambda:[0,T]\times\Omega\times A\to{\mathbb{R}}_{+}$, $\ell:[0,T]\times\Omega\times A\to{\mathbb{R}}$, and $h:\Omega\to{\mathbb{R}}$.
Assumption 7.1.
Suppose that $f$, $\lambda$, $\ell$, and $h$ satisfy the following conditions.
(i) For a.e. $t\in[0,T]$, the map $(x,y,z)\mapsto(f,\lambda,\ell)(x,a)$, $\Omega\times A\to{\mathbb{R}}^{d}\times{\mathbb{R}}_{+}\times{\mathbb{R}}$, is continuous. Furthermore, $h$ is continuous.
(ii) For every $(x,a)\in\Omega\times A$, the map $t\mapsto(f,\lambda,\ell)(t,x,a)$, $[0,T]\to{\mathbb{R}}^{d}\times{\mathbb{R}}_{+}\times{\mathbb{R}}$, is Borel measurable.
(iii) There are constants $C_{f}$, $C_{\lambda}\geq 0$ such that, for a.e. $t\in(0,T)$ and each $x\in\Omega$,
$\sup_{a\in A}\left(\left|{f(t,x,a))}\right|+\left|{\ell(t,x,a)}\right|\right)\leq C_{f}(1+\sup_{s\leq t}\left|{x(s)}\right|)\text{ and }\sup_{a\in A}\left|{\lambda(t,x,a)}\right|\leq C_{\lambda}.$
(iv) There is a constant $L_{f}\geq 0$ such that, for a.e. $t\in(0,T)$ and every $x$, $\tilde{x}\in\Omega$,
$\sup_{a\in A}\left(\left|{f(t,x,a)-f(t,\tilde{x},a))}\right|+\left|{\ell(t,x,a)-\ell(t,\tilde{x},a))}\right|+\left|{\lambda(t,x,a)-\lambda(t,\tilde{x},a)}\right|\right)$
$\qquad\leq L_{f}\,\sup_{s\leq t}\left|{x(s)-\tilde{x}(s)}\right|.$
For each $(t_{0},x_{0},\alpha)\in[0,T)\times\Omega\times\mathcal{A}$, let $\phi^{t_{0},x_{0},\alpha}=\phi$ be the solution of
$\phi^{\prime}(t)$ $=f(t,\phi,\alpha(t))\text{ a.e.~{}on $(t_{0},T)$},\quad\phi(t)=x_{0}(t)\text{ on $[0,t_{0}]$}.$
Moreover, define, for each $(s,x,a)\in[0,T)\times\Omega\times\mathcal{A}$ and $t\in[s,T]$,
$\lambda^{s,x,a}(t)$ $:=\lambda(t,\phi^{s,x,a},a(t)),$
$\chi^{s,x,a}(t)$ $:=\exp\left(-\int_{s}^{t}\lambda^{s,x,a}(r)\,{\mathrm{d}}r\right),$
$\ell^{s,x,a}(t)$ $:=\ell(t,\phi^{s,x,a},a(t)).$
Our goal is to show that the value function $v:[0,T]\times\Omega\to{\mathbb{R}}$ defined by
(7.1) $v(s,x):=\inf_{a\in\mathcal{A}}\left[\int_{s}^{T}\chi^{s,x,a}(t)\,\ell^{s,x,a}(t)\,{\mathrm{d}}t+\chi^{s,x,a}(T)\,h(\phi^{s,x,a})\right]$
is a minimax solution of
(7.2) $\begin{split}-\partial_{t}u(t,x)-H(t,x,u(t,x),\partial_{x}u(t,x))&=0,\quad(t,x)\in[0,T)\times\Omega,\\ u(T,x)&=h(x),\quad x\in\Omega,\end{split}$
where $H:[0,T]\times\Omega\times{\mathbb{R}}\times{\mathbb{R}}^{d}\to{\mathbb{R}}$ is defined by
(7.3) $H(t,x,y,z):=\operatornamewithlimits{ess\,inf}_{a\in A}\left[\ell(t,x,a)+(f(t,x,a),z)-\lambda(t,x,a)\,y\right].$
Remark 7.2.
Suppose that Assumption 7.1 holds. One can show that $H$ defined by (7.3) satisfies Assumption 4.2 with $L_{H}=C_{f}$ and $M_{L}$ independent from $L$.
The dynamic programming principle holds. Its proof is standard and is omitted.
Proposition 7.3.
If $0\leq s\leq t\leq T$ and $x\in\Omega$, then
$v(s,x)=\inf_{a\in\mathcal{A}}\left[\int_{s}^{t}\chi^{s,x,a}(r)\,\ell^{s,x,a}(r)\,{\mathrm{d}}r+\chi^{s,x,a}(t)\,v(t,\phi^{s,x,a})\right].$
Proposition 7.4.
For each $L\geq 0$ and $x_{0}\in\Omega$, there is a $C_{L,x_{0}}\geq 0$, such that $0\leq t_{0}\leq t_{1}\leq T$ and $x$, $\tilde{x}\in\mathcal{X}^{L}(t_{0},x_{0})$ imply $\left|{v(t_{0},x)-v(t_{1},x)}\right|\leq C_{L,x_{0}}(t_{1}-t_{0})$ and $\left|{v(t_{0},x)-v(t_{0},\tilde{x})}\right|\leq C_{L,x_{0}}\sup_{s\leq t_{0}}\left|{x(s)-\tilde{x}(s)}\right|$. Moreover, $v$ is continuous.
Proof.
Proceed as in the proof of Theorem 7.4 in [BK18JFA]. ∎
Lemma 7.5.
Let $u$ be a minimax $L_{f}$-subsolution. Let $v$ be the value function defined by (7.1). Then the function $(t,x,\tilde{x})\mapsto w(t,x,\tilde{x}):=u(t,x)-v(t,\tilde{x})$, $[0,T]\times\Omega\times\Omega\to{\mathbb{R}}$, is a viscosity $C_{f}$-subsolution of (5.2) with parameter $\Upsilon=u$.
Proof.
Fix $(s_{0},x_{0},\tilde{x}_{0})\in[0,T)\times\Omega\times\Omega$ and $\varphi\in\underline{\mathcal{A}}^{(C_{f})}w(s_{0},x_{0},\tilde{x}_{0})$ with corresponding time $T_{0}\in(s_{0},T]$. Let $w(s_{0},x_{0},\tilde{x}_{0})>0$. Put
$y_{0}:=u(s_{0},x_{0}),\,\tilde{y}_{0}:=v(s_{0},x_{0}),\,z:=\partial_{x}\varphi(s_{0},x_{0},\tilde{x}_{0}),\,\tilde{z}:=-\partial_{\tilde{x}}\varphi(s_{0},x_{0},\tilde{x}_{0}).$
Also, for every $a\in\mathcal{A}$, put
$\phi^{a}:=\phi^{s_{0},\tilde{x}_{0},a},\,\lambda^{a}:=\lambda^{s_{0},\tilde{x}_{0},a},\,\chi^{a}:=\chi^{s_{0},\tilde{x}_{0},a},\,\ell^{a}:=\ell^{s_{0},\tilde{x}_{0},a},\,f^{a}:=f(\cdot,\phi^{a},a(\cdot)).$
First note that, due to the minimax $C_{f}$-subsolution property of $u$, we can fix a pair $(x,y)\in\mathcal{Y}^{C_{f}}(s_{0},x_{0},y_{0},z)$ such that, for all $t\in[s_{0},T]$, we have
(7.4) $u(t,x)-y_{0}\geq y(t)-y_{0}=\int_{s_{0}}^{t}\left[(x^{\prime}(s),z)-H(s,x,y(s),z)\right]\,{\mathrm{d}}s.$
Next, let $\delta\in(0,T_{0}-s_{0}]$ and let ${\varepsilon}>0$. By Proposition 7.3, there exists a control $a=a^{\delta,{\varepsilon}}\in\mathcal{A}$ such that
$\tilde{y}_{0}=v(s_{0},x_{0})>\chi^{a}(s_{0}+\delta)\,v(s_{0}+\delta,\phi^{a})+\int_{s_{0}}^{s_{0}+\delta}\chi^{a}(s)\,\ell^{a}(s)\,{\mathrm{d}}s-\delta\,{\varepsilon}.$
Thus
(7.5) $\begin{split}&-\left[v(s_{0}+\delta,\phi^{a})-\tilde{y}_{0}\right]\\ &\qquad>-\delta\,{\varepsilon}+[\chi^{a}(s_{0}+\delta)-1]\,v(s_{0}+\delta,\phi^{a})+\int_{s_{0}}^{s_{0}+\delta}\chi^{a}(s)\,\ell^{a}(s)\,{\mathrm{d}}s\\ &\qquad=-\delta\,{\varepsilon}+\int_{s_{0}}^{s_{0}+\delta}\Bigl{\{}[-\lambda^{a}(s)]\,\chi^{a}(s)\,v(s_{0}+\delta,\phi^{a})+\chi^{a}(s)\,\ell^{a}(s)\Bigr{\}}\,{\mathrm{d}}s.\end{split}$
Now we use the test function property of $\varphi$, i.e., we have $(\varphi-w)(s_{0},x_{0},\tilde{x}_{0})\leq(\varphi-w)(s_{0}+\delta,x,\phi^{a})$, which together with (7.4) and (7.5) yield
$\int_{s_{0}}^{s_{0}+\delta}\left[\partial_{t}\varphi(s,x,\phi^{a})+(x^{\prime}(s),\partial_{x}\varphi(s,x,\phi^{a}))+(f^{a}(s),\partial_{\tilde{x}}\varphi(s,x,\phi^{a}))\right]\,{\mathrm{d}}s$
$=\varphi(s_{0}+\delta,x,\phi^{a})-\varphi(s_{0},x_{0},\tilde{x}_{0})$
$\geq\left[u(s_{0}+\delta,x)-v(s_{0}+\delta,\phi^{a})\right]-\left[y_{0}-\tilde{y}_{0}\right]$
$>\int_{s_{0}}^{s_{0}+\delta}\left[(x^{\prime}(s),z)-H(s,x,y(s),z)-\lambda^{a}(s)\,\chi^{a}(s)\,v(s_{0}+\delta,\phi^{a})+\chi^{a}(s)\,\ell^{a}(s)\right]\,{\mathrm{d}}s$
$\qquad-\delta\,{\varepsilon}.$
Consequently,
$-\delta\,{\varepsilon}$ $<\int_{s_{0}}^{s_{0}+\delta}\Bigl{\{}\partial_{t}\varphi(s,x,\phi^{a})+H(s,x,y(s),z)-\left[\ell^{a}(s)+(f^{a}(s),\tilde{z})-\lambda^{a}(s)\,\tilde{y}_{0}\right]$
$\qquad\qquad\quad+(x^{\prime}(s),\partial_{x}\varphi(s,x,\phi^{a})-z)+(f^{a}(s),\tilde{z}+\partial_{\tilde{x}}\varphi(s,x,\phi^{a}))$
$\qquad\qquad\quad+[1-\chi^{a}(s)]\,\ell^{a}(s)+\lambda^{a}(s)\,\chi^{a}(s)\,v(s_{0}+\delta,\phi^{a})-\lambda^{a}(s)\,\tilde{y}_{0}\Bigr{\}}\,{\mathrm{d}}s$
$\leq\int_{s_{0}}^{s_{0}+\delta}\Bigl{\{}\partial_{t}\varphi(s,x,\phi^{a})+H(s,x,y(s),z)-H(s,\phi^{a},\tilde{y}_{0},\tilde{z})$
$\qquad\qquad\quad+(x^{\prime}(s),\partial_{x}\varphi(s,x,\phi^{a})-z)+(f^{a}(s),\tilde{z}+\partial_{\tilde{x}}\varphi(s,x,\phi^{a}))\Bigr{\}}\,{\mathrm{d}}s$
$\qquad+I^{a}(s_{0}+\delta)+J^{a}(s_{0}+\delta),$
where
$I^{a}(s_{0}+\delta)$ $:=\int_{s_{0}}^{s_{0}+\delta}[1-\chi^{a}(s)]\,\ell^{a}(s)\,{\mathrm{d}}s=\int_{s_{0}}^{s_{0}+\delta}\left[\int_{s_{0}}^{s}\lambda^{a}(r)\,\chi^{a}(r)\,{\mathrm{d}}r\right]\,\ell^{a}(s)\,{\mathrm{d}}s,$
$J^{a}(s_{0}+\delta)$ $:=\int_{s_{0}}^{s_{0}+\delta}\left[\lambda^{a}(s)\,\chi^{a}(s)\,v(s_{0}+\delta,\phi^{a})-\lambda^{a}(s)\,\tilde{y}_{0}\right]\,{\mathrm{d}}s,$
$=\int_{s_{0}}^{s_{0}+\delta}\Bigl{\{}\lambda^{a}(s)\,[\chi^{a}(s)-1]\,v(t,\phi^{a})+\lambda^{a}(s)\,v(s_{0}+\delta,\phi^{a})-\lambda^{a}(s)\,\tilde{y}_{0}\Bigr{\}}\,{\mathrm{d}}s$
$=J^{a}_{1}(s_{0}+\delta)+J^{a}_{2}(s_{0}+\delta),$
$J^{a}_{1}(s_{0}+\delta)$ $:=\int_{s_{0}}^{s_{0}+\delta}\lambda^{a}(s)\,\int_{s_{0}}^{s}[-\lambda^{a}(r)\,\chi^{a}(r)]\,{\mathrm{d}}r\,{\mathrm{d}}s\cdot v(s_{0}+\delta,\phi^{a}),\text{ and}$
$J^{a}_{2}(s_{0}+\delta)$ $:=\int_{s_{0}}^{s_{0}+\delta}\lambda^{a}(s)\,{\mathrm{d}}s\cdot[v(s_{0}+\delta,\phi^{a})-\tilde{y}_{0}].$
We estimate now those error terms. By Assumption 7.1,
$\left|{I^{a}(s_{0}+\delta)}\right|$ $\leq C_{\lambda}\,C_{f}\,(1+\sup\nolimits_{s\leq s_{0}+\delta}\left|{\phi^{a}(s)}\right|)\,\delta^{2}\leq C_{1}\,\delta^{2}$
for some constant $C_{1}$ that depends only on $\tilde{x}_{0}$ because $\phi^{a}\in\mathcal{X}^{(C_{f})}(s_{0},\tilde{x}_{0})$. By Assumption 7.1 and Proposition 7.4,
$\left|{J^{a}_{1}(s_{0}+\delta)}\right|$ $\leq C_{\lambda}^{2}\,\delta^{2}\,\left|{v(s_{0}+\delta,\phi^{a})}\right|\leq C_{2}\,\delta^{2}\text{ and}$
$\left|{J^{a}_{2}(s_{0},\delta)}\right|$ $\leq(C_{\lambda}\,\delta)\,\left|{v(s_{0}+\delta,\phi^{a})-v(s_{0},\phi^{a})}\right|\leq C_{2}\,\delta^{2}$
for some constant $C_{2}$ that depends only on $\tilde{x}_{0}$. Therefore, for all sufficiently large $n\in{\mathbb{N}}$, we have, with $\tilde{x}^{n}:=\phi^{a^{\delta,{\varepsilon}}}|_{\delta={\varepsilon}=n^{-1}}$,
(7.6) $\begin{split}&-\frac{1}{n^{2}}-\frac{C_{1}+2C_{2}}{n^{2}}\\ &\qquad\leq\int_{s_{0}}^{s_{0}+n^{-1}}\Bigl{[}\partial_{t}\varphi(s,x,\tilde{x}^{n})+H(s,x,y(s),z)-H(s,\tilde{x}^{n},\tilde{y}_{0},\tilde{z})\\ &\qquad\qquad+(x^{\prime}(s),\partial_{x}\varphi(s,x,\tilde{x}^{n})-z)+((\tilde{x}^{n})^{\prime}(s),\tilde{z}+\partial_{\tilde{x}}\varphi(s,x,\tilde{x}^{n}))\Bigr{]}\,{\mathrm{d}}s\\ &\qquad\leq\int_{s_{0}}^{s_{0}+n^{-1}}\Bigl{[}\partial_{t}\varphi(s,x,\tilde{x}^{n})+M_{L}(1+\left|{y(s)}\right|+\left|{z}\right|)\sup_{t\leq s}\left|{x(t)-\tilde{x}^{n}(t)}\right|\\ &\qquad\qquad+L_{H}(1+\sup_{t\leq s}\left|{\tilde{x}^{n}(t)}\right|)\left|{z-\tilde{z}}\right|+0\\ &\qquad\qquad+(x^{\prime}(s),\partial_{x}\varphi(s,x,\tilde{x})-z)+((\tilde{x}^{n})^{\prime}(s),\partial_{\tilde{x}}\varphi(s,x,\tilde{x}^{n})+\tilde{z})\Bigr{]}\,{\mathrm{d}}s,\end{split}$
where the last inequality can be derived exactly as in (5.5). Finally, let $\tilde{x}$ be a limit in $\mathcal{X}^{(C_{f})}(s_{0},\tilde{x}_{0})$ of a convergent subsequence $(\tilde{x}^{n_{k}})_{k}$ of $(\tilde{x}^{n})_{n}$ such that also $((\tilde{x}^{n_{k}})^{\prime})_{k}$ converges weakly to $\tilde{x}^{\prime}$ in $L^{2}(s_{0},T;{\mathbb{R}}^{d})$. Replacing $n$ by $n_{k}$ in (7.6) and letting $k\to\infty$ yields (5.3). This concludes the proof. ∎
Lemma 7.6.
Let $v$ be the value function defined by (7.1). Let $u$ be a minimax $C_{f}$-supersolution. Then the function $(t,x,\tilde{x})\mapsto w(t,x,\tilde{x}):=v(t,x)-u(t,\tilde{x})$, $[0,T]\times\Omega\times\Omega\to{\mathbb{R}}$, is a viscosity $C_{f}$-subsolution of (5.2) with parameter $\Upsilon=v$.
Proof.
Let $(s_{0},x_{0},\tilde{x}_{0})\in[0,T)\times\Omega\times\Omega$. Let $\varphi\in\underline{\mathcal{A}}^{(C_{f})}w(s_{0},x_{0},\tilde{x}_{0})$ with corresponding time $T_{0}\in(s_{0},T]$. Suppose that $w(s_{0},x_{0},\tilde{x}_{0})>0$. Put
$y_{0}$ $:=v(s_{0},x_{0}),\quad\tilde{y}_{0}:=u(s_{0},\tilde{x}_{0}),\quad z:=\partial_{x}\varphi(s_{0},x_{0},\tilde{x}_{0}),\quad\tilde{z}:=-\partial_{\tilde{x}}\varphi(s_{0},x_{0},\tilde{x}_{0}).$
For every $a\in\mathcal{A}$, set
$\phi^{a}$ $:=\phi^{s_{0},x_{0},a},\,\lambda^{a}:=\lambda^{s_{0},x_{0},a},\,\chi^{a}:=\chi^{s_{0},x_{0},a},\,\ell^{a}:=\ell^{s_{0},x_{0},a},\,f^{a}:=f(\cdot,\phi^{a},a(\cdot)).$
Since $u$ is a minimax $C_{f}$-supersolution, there is a pair $(\tilde{x},\tilde{y})\in\mathcal{Y}^{(C_{f})}(s_{0},\tilde{x}_{0},\tilde{y}_{0},\tilde{z})$ such that, for all $t\in[s_{0},T]$, we have
(7.7) $u(t,\tilde{x})-\tilde{y}_{0}\leq\tilde{y}(t)-\tilde{y}_{0}=\int_{s_{0}}^{t}\left[(\tilde{x}^{\prime}(s),\tilde{z})-H(s,\tilde{x},\tilde{y}(s),\tilde{z})\right]\,{\mathrm{d}}s.$
Next, let $\delta\in(0,T_{0}-s_{0}]$ and $a\in\mathcal{A}$. By Proposition 7.3,
$y_{0}=v(s_{0},x_{0})\leq\chi^{a}(s_{0}+\delta)\,v(s_{0}+\delta,\phi^{a})+\int_{s_{0}}^{s_{0}+\delta}\chi^{a}(s)\,\ell^{a}(s)\,{\mathrm{d}}s.$
Thus
(7.8) $\begin{split}v(s_{0}+\delta,\phi^{a})-y_{0}&\geq[1-\chi^{a}(s_{0}+\delta)]\,v(s_{0}+\delta,\phi^{a})-\int_{s_{0}}^{s_{0}+\delta}\chi^{a}(s)\,\ell^{a}(s)\,{\mathrm{d}}s\\ &=\int_{s_{0}}^{s_{0}+\delta}\left[\lambda^{a}(s)\,\chi^{a}(s)\,v(s_{0}+\delta,\phi^{a})-\chi^{a}(s)\,\ell^{a}(s)\right]\,{\mathrm{d}}s.\end{split}$
By (7.7), (7.8), and the test function property of $\varphi$,
$\int_{s_{0}}^{s_{0}+\delta}\left[\partial_{t}\varphi(s,\phi^{a},\tilde{x})+(f^{a}(s),\partial_{x}\varphi(s,\phi^{a},\tilde{x}))+(\tilde{x}^{\prime}(s),\partial_{\tilde{x}}\varphi(s,\phi^{a},\tilde{x}))\right]\,{\mathrm{d}}s$
$\quad\geq[v(s_{0}+\delta,\phi^{a})-u(s_{0}+\delta,\tilde{x})]-[y_{0}-\tilde{y}_{0}]$
$\quad\geq\int_{s_{0}}^{s_{0}+\delta}\left[\lambda^{a}(s)\,\chi^{a}(s)\,v(s_{0}+\delta,\phi^{a})-\chi^{a}(s)\,\ell^{a}(s)-(\tilde{x}^{\prime}(s),\tilde{z})+H(s,\tilde{x},\tilde{y}(s),\tilde{z})\right]\,{\mathrm{d}}s.$
Therefore,
(7.9) $\begin{split}&\int_{s_{0}}^{s_{0}+\delta}\Bigl{\{}\partial_{t}\varphi(s,\phi^{a},\tilde{x})+[\ell^{a}(s)+(f^{a}(s),z)-\lambda^{a}(s)\,y_{0}]-H(s,\tilde{x},\tilde{y}(s),\tilde{z})\\ &\qquad\qquad+(f^{a}(s),\partial_{x}\varphi(s,\phi^{a},\tilde{x})-z)+(\tilde{x}^{\prime}(s),\partial_{\tilde{x}}\varphi(s,\phi^{a},\tilde{x})+\tilde{z})\Bigr{\}}\,{\mathrm{d}}s\\ &\quad\geq\int_{s_{0}}^{s_{0}+\delta}\Bigl{\{}\lambda^{a}(s)\,\chi^{a}(s)\,v(s_{0}+\delta,\phi^{a})-\lambda^{a}(s)\,y_{0}+(1-\chi^{a})\,\ell^{a}(s)\Bigr{\}}\,{\mathrm{d}}s\\ &\quad\geq-C\delta^{2}\end{split}$
for some constant $C>0$ independent from $\delta$ and $a$. The last inequality in (7.9) can be shown exactly as the estimation of the terms $I^{a}(s_{0}+\delta)$ and $J^{a}(s_{0}+\delta)$ in the proof of Lemma 7.5.
We continue now by proceeding similarly as in the Step 1 of proof of Theorem 6.7 in [Qiu18SICON_SHJB], i.e., we obtain, via a measurable selection argument (e.g., Theorem 21.3.4 in [CohenElliott]), for every ${\varepsilon}>0$ the existence of a control $a^{\varepsilon}\in\mathcal{A}$ such that
$\ell^{a^{\varepsilon}}(s)+(f^{a^{\varepsilon}}(s),z)-\lambda^{a^{\varepsilon}}(s)\,y_{0}\leq H(s,\phi^{a^{\varepsilon}},y_{0},z)+{\varepsilon}\quad\text{a.e.~{}in $(s_{0},T_{0})$.}$
Thus, together with (7.9),
(7.10) $\begin{split}&-C\delta^{2}-\delta{\varepsilon}\leq\int_{s_{0}}^{s_{0}+\delta}\Bigl{[}\partial_{t}\varphi(s,\phi^{a^{\varepsilon}},\tilde{x})+H(s,\phi^{a^{\varepsilon}},y_{0},z)-H(s,\tilde{x},\tilde{y}(s),z)\\ &\qquad\qquad\qquad+(f^{a^{\varepsilon}}(s),\partial_{x}\varphi(s,\phi^{a^{\varepsilon}},\tilde{x})-z)+(\tilde{x}^{\prime}(s),\partial_{\tilde{x}}\varphi(s,\phi^{a^{\varepsilon}},\tilde{x})+\tilde{z})\Bigr{]}\,{\mathrm{d}}s\\ &\qquad\leq\int_{s_{0}}^{s_{0}+\delta}\Bigl{[}\partial_{t}\varphi(s,\phi^{a^{\varepsilon}},\tilde{x})+M_{L}(1+\left|{y_{0}}\right|+\left|{z}\right|)\sup_{t\leq s}\left|{\phi^{a^{{\varepsilon}}}(t)-\tilde{x}(t)}\right|\\ &\qquad\qquad\qquad+L_{H}(1+\sup_{t\leq s}\left|{\tilde{x}(t)}\right|)\left|{z-\tilde{z}}\right|\\ &\qquad\qquad\qquad+(f^{a^{\varepsilon}}(s),\partial_{x}\varphi(s,\phi^{a^{\varepsilon}},\tilde{x})-z)+(\tilde{x}^{\prime}(s),\partial_{\tilde{x}}\varphi(s,\phi^{a^{\varepsilon}},\tilde{x})+\tilde{z})\Bigr{]}\,{\mathrm{d}}s.\end{split}$
We refer to (5.5) for details regarding the last inequality. Finally, dividing (7.10) by $\delta$, letting $\delta\downarrow 0$, and noting that ${\varepsilon}>0$ was arbitrary yields (5.3), which concludes the proof. ∎
Theorem 7.7.
The value function $v$ is the unique minimax $C_{f}$-solution of (7.2).
Proof.
By Theorem 6.4, there is a unique minimax $C_{f}$-solution $u$ of of (7.2). By Theorem 5.3 and Lemma 7.5, $u\leq v$. By Theorem 5.3 and Lemma 7.6, $v\leq u$. ∎